What is the inverse function of f(x)=x^2+x, x ≥ -1/2?
Found 2 solutions by Theo, solver91311:
Answer by Theo(13342)

actually, x needs to be greater than or equal to -1/4, not -1/2, as you will see later.

if you let y = f(x), then your function of f(x) = x^2 + x becomes:

y = x^2 + x

since x^2 + x = (x + 1/2)^2 - 1/4, then your function becomes:

y = (x + 1/2)^2 - 1/4

add 1/4 to both sides of this function to get:

y + 1/4 = (x + 1/2)^2

take the square root of both sides of this function to get:

sqrt(y + 1/4) = x + 1/2

subtract 1/2 from both sides of the function to get:

sqrt(y + 1/4) - 1/2 = x

replace x with y and y with x to get:

sqrt(x + 1/4) - 1/2 = y

switch sides to get:

y = sqrt(x + 1/4) - 1/2

that's your inverse function.

if it is truly the inverse function, then it must be a reflection about the line y = x and (x,y) in the normal function must be the same distance from the line y = x as (y,x) is in the inverse function.

i graphed both the normal function and the inverse function and the line y = x to show you that the inverse function is a reflection about the line y = x.

i also graphed the line y = -x + 6 to show you that the point (x,y) in the normal function is the same distance from the line y = x as the point (y,x) in the inverse function.

here's what the graph looks like.



the distance between the point (1.646,4.354) on the normal graph and the point (3,3) on the line y = x is the same distance between the point (4.354,1.646) on the inverse graph and the same point of (3,3) on the line y = x.

to confirm, use the distance formula of d = sqrt((x2-x1)^2 + (y2-y1)^2)

when (x1,y1) from the normal function = (1.646,4.354) and (x2,y2) from the line (y = x) = (3,3), then:

d = sqrt((3 - 1.646)^2 + (3 - 4.354)^2) = 1.914845163

when (x1,y1) from the inverse function = (4.354,1.646)) and (x2,y2) from the line (y = x) = (3,3), then:

d = sqrt(3 - 4.354)^2 + (3 - 1.646)^2) = 1.914845163

the distances are the same, confirming that the inverse function for y = x^2 + x is the function y = sqrt(x + 1/4) - 1/2.



Answer by solver91311(24713)





The domain is (given) and the range is



Complete the square



Factor the RHS



Take the square root



Isolate



Swap the variables and swap the domain and range:



The domain is and the range is

Eliminating the negative side of the square root based on the range restriction:



John

My calculator said it, I believe it, that settles it


I > Ø